Nicole:

On a vacation, DeVante and his family traveled 385 miles at 65 mph. How long did the trip take? 5.96 hours 5.59 hours 5.69 hours 5.92 hours

Nicole:

@TheSmartOne

TheSmartOne:

We have miles and we have the speed in miles per hour How would you set it up so that the miles can cross out and all you're left with is hours in the numerator

Nicole:

65/385x ?

TheSmartOne:

Err not quite It might help you to keep the units 385 miles * 65 miles/hour OR 385 miles * 1 hour/65 miles Which one would it be?

Nicole:

No

Nicole:

The second one

Nicole:

Like 385/x=65/1

TheSmartOne:

Good sav \( \sf385~miles \times \frac{1~hour}{65~miles}\)

TheSmartOne:

so miles can cross out and you can do the simple division to get the time it took for them to travel

Nicole:

Which is 5.92

TheSmartOne:

Spot on!

jhonyy9:

the other easy way is the rule of 3 simple for the same problems given 65 miles in 1 hour 385 miles in x h cross multiply and calcule x x = 385*1 / 65 = ?

Nicole:

yes 5.92

Nicole:

thank you

jhonyy9:

np my pleasure